OpenStudy (anonymous):

if you were to travel to a star 50 light years from earth at a speed of 2.0 X 10 to 8th power m/s, what would you measure this distance to be?

OpenStudy (calculusfunctions):

@shortcake1970 do you know the definition of a light year?

OpenStudy (calculusfunctions):

A light year is the distance that light travels in one Earth year.

OpenStudy (anonymous):

light travels at 3*10^8 m/s

OpenStudy (calculusfunctions):

Yes that's approximately the speed of light, as @VeritasVosLiberabit said.

OpenStudy (anonymous):

so it seems the 2*10^8 is meant to throw you off because it is irrelevant to finding the distance

OpenStudy (anonymous):

and the distance light travels in a year is 9.47 X 10^15 i have that much

OpenStudy (calculusfunctions):

Now find the number of seconds in one Earth year. Yes @VeritasVosLiberabit , that is extraneous information. Not needed.

OpenStudy (anonymous):

it might be better to keep your calculations in MKS units unless told otherwise.

OpenStudy (calculusfunctions):

Alright @shortcake1970 now simply multiply that answer by 50.

OpenStudy (anonymous):

so do I just multiply?

OpenStudy (calculusfunctions):

Haha that's what I said.

OpenStudy (anonymous):

oops yup you just answered that

OpenStudy (anonymous):

Thank you!

OpenStudy (calculusfunctions):

Welcome!

OpenStudy (calculusfunctions):

Need help with anything else or are you okay?

OpenStudy (anonymous):

well I do have another questions I might need help with let me write it...

OpenStudy (calculusfunctions):

Sure

OpenStudy (calculusfunctions):

Are these questions you've tried yourself? It's important you try them yourself first.

OpenStudy (anonymous):

a beam of a certain type of elementary particle travels at a speed of 2.8 X 10^8 m/s. at this speed the average lifetime is measured to be 2.5 X 10^-8 s. what is the particle's lifetime at rest. let me tell you what formula i was thinking of using

OpenStudy (anonymous):

once i know which formula i can work it out myself :)

OpenStudy (calculusfunctions):

Sorry I was away for a minute. Just give me second. I'm just in the middle of something.

OpenStudy (anonymous):

ok i'm having a hard time typing the formula

OpenStudy (anonymous):

delta t=delta t o /square root of 1-v^2/c^2

OpenStudy (calculusfunctions):

OK I'm back. Yes\[\Delta t =\frac{ \Delta t _{0} }{ \sqrt{1-\frac{ v ^{2} }{ c ^{2} }} }\]

OpenStudy (calculusfunctions):

Do you understand? \[\Delta t = 2.5\times10^{-8}s\]\[v =2.8\times10^{8}m/s\]\[c =3.0\times10^{8}m/s\]\[\Delta t _{0}=?\]Rearrange the formula first to get\[\Delta t _{0}=\Delta t \sqrt{1-\frac{ v ^{2} }{ c ^{2} }}\]

OpenStudy (anonymous):

.90 x 10^8 s

OpenStudy (calculusfunctions):

Give me a second to calculate it.

OpenStudy (anonymous):

is the answer for the previous question 4.73 x 10^17?

OpenStudy (calculusfunctions):

It should be\[9.0\times10^{-9}\]

OpenStudy (calculusfunctions):

That exponent is negative 9

OpenStudy (anonymous):

oh i see

OpenStudy (calculusfunctions):

Do you? Did you spot your error?

OpenStudy (anonymous):

yes i spotted the error

OpenStudy (calculusfunctions):

Great. I have to go but is there anything else before I do?

OpenStudy (calculusfunctions):

I guess not. I might be back a little later. Bye.

OpenStudy (anonymous):

thank you so much.